# Deep Learning for Protein-Ligand Docking: Are We There Yet?

### v5

@@ Abstract
The effects of ligand binding on protein structures and their in vivo functions carry numerous implications for modern biomedical research and biotechnology development efforts such as drug discovery. Although several deep learning (DL) methods and benchmarks designed for protein-ligand docking have recently been introduced, to date no prior works have systematically studied the behavior of the latest docking and structure prediction methods within the broadly applicable context of (1) using predicted (apo) protein structures for docking (e.g., for applicability to new proteins); (2) binding multiple (cofactor) ligands concurrently to a given target protein (e.g., for enzyme design); and (3) having no prior knowledge of binding pockets (e.g., for generalization to unknown pockets). To enable a deeper understanding of docking methods’ real-world utility, we introduce PoseBench, the first comprehensive benchmark for broadly applicable protein-ligand docking. PoseBench enables researchers to rigorously and systematically evaluate DL methods for apo-to-holo protein-ligand docking and protein-ligand structure prediction using both primary ligand and multi-ligand benchmark datasets, the latter of which we introduce for the first time to the DL community. Empirically, using PoseBench, we find that (1) DL co-folding methods generally outperform comparable conventional and DL docking baselines, yet popular methods such as AlphaFold 3 are still challenged by prediction targets with novel protein sequences; (2) certain DL co-folding methods are highly sensitive to their input multiple sequence alignments, while others are not; and (3) DL methods struggle to strike a balance between structural accuracy and chemical specificity when predicting novel or multi-ligand protein targets. Code, data, tutorials, and benchmark results are available at https://github.com/BioinfoMachineLearning/PoseBench.
@@ 1 Introduction
As the field has gradually begun to investigate whether proteins in complex with other types of molecules can faithfully be modeled with ML (and particularly deep learning (DL)) techniques [5, 6, 7], several new works in this direction have suggested the promising potential of such approaches to protein-ligand structure determination [8, 9, 10, 11]. Nonetheless, it remains to be shown the extent to which the latest of such DL methods can adequately generalize to the context of binding novel or uncommon protein-ligand interaction (PLI) pockets and multiple interacting ligand molecules (e.g., which can alter the chemical functions of various enzymes) as well as whether such methods can faithfully model amino acid-specific types of PLIs natively found in crystallized biomolecular structures.
Figure 1: Overview of PoseBench, our comprehensive benchmark for broadly applicable DL modeling of primary and multi-ligand protein complex structures. Baseline methods of the benchmark include a range of the latest DL docking and co-folding methods, both open-source and commercially restrictive, as well as conventional algorithms for docking. Key observations derived using PoseBench include the discontinuity between structure and interaction modeling performance for novel or uncommon prediction targets and the heavy reliance of key DL co-folding methods on MSA-based input features to achieve high structural accuracy.
@@
  • We introduce the first unified benchmark for protein-ligand docking and structure prediction that evaluates the performance of several recent DL-based methods (e.g., AlphaFold 3, Chai-1) as well as conventional algorithms (e.g., AutoDock Vina) for primary and multi-ligand docking, which suggests that DL co-folding methods generally outperform conventional algorithms yet remain challenged by novel or uncommon prediction targets.
  • In contrast to several recent works using crystal protein structures for protein-ligand docking [12, 13], the docking benchmark results we present in this work are all within the context of standardized input multiple sequence alignments (MSAs) and high accuracy apo-like (i.e., AlphaFold 3-predicted) protein structures without specifying known binding pockets, which notably enhances the broad applicability of this study’s findings.
  • Our newly proposed benchmark, PoseBench, enables specific insights into necessary areas of future work for accurate and generalizable biomolecular structure prediction, including that DL methods struggle to balance faithful modeling of native PLI fingerprints (PLIFs) with structural accuracy during pose prediction and that some DL co-folding methods are more dependent than others on the availability of input MSAs.
  • Our benchmark results also highlight the importance of including challenging (out-of-sequence-distribution) datasets when evaluating future DL methods while measuring their ability to recapitulate amino acid-specific PLIFs with an appropriate new metric that we introduce in this work.
@@ 2.1 Astex Diverse results
Containing PLI structures deposited in the RCSB Protein Data Bank (PDB) [15] up until 2007, most of the well-known Astex Diverse dataset’s structures [16] are present in the training data of each baseline method, yet benchmarking results for this dataset (n=85), shown in Figure 2, indicate that only DL co-folding methods achieve higher structural and chemical accuracy rates (RMSD $\leq$ 2 Å & PB-Valid) than the conventional docking baseline AutoDock Vina combined with P2Rank for PLI binding site prediction to facilitate blind molecular docking. Interestingly, nearly all baseline methods identify the correct PLI binding pocket approximately 90% of the time, yet only the DL co-folding methods AlphaFold 3 (AF3) [11] and Chai-1 [17] achieve a reasonable balance between their rates of structural and chemical accuracy and chemical specificity (PLIF-WM), with the single-sequence (i.e., MSA-ablated) version of AF3 being a notable exception. These results suggest that DL co-folding methods have learned the most comprehensive representations of this dataset’s input sequences, yet only the performance of the DL co-folding method Chai-1 is maintained without the availability of diverse input MSAs. One likely explanation for this phenomenon is that Chai-1’s training primarily relied upon the availability of amino acid sequence embeddings generated by the protein language model ESM2 [18] in addition to features derived from input MSAs, which may have imbued the model with rich MSA-independent representations for biomolecular structure prediction.
@@ 2.2 DockGen-E results
As visualized in Figure 3, results with our new DockGen-E dataset of biologically relevant PLI complexes deposited in the PDB up to 2019 (n=122) demonstrate that only the latest DL co-folding methods can locate a sizable fraction of structurally accurate PLI binding poses represented in this dataset. As such methods may have previously seen these PLI structures in their respective training data, it is surprising that even the latest AF3 model fails to identify a structurally and chemically accurate pose for more than half of the dataset’s complexes. Further, for Chai-1 and AF3, their single-sequence variants achieve slightly higher chemical specificity than their MSA-based versions, which may indicate that for these methods MSA features obfuscate primary sequence knowledge in favor of evolution-averaged (i.e., amino acid-generic) representations. The overall lower range of PLIF-WM values achieved by each method for this dataset further suggests the increased chemical modeling difficulty of this dataset’s complexes compared to those presented by the Astex Diverse dataset. A potential source of these difficulties is that each of this dataset’s complexes represents a functionally distinct PLI binding pocket (as codified by ECOD domains [19], see [20] for more details) compared to data deposited in the PDB before 2019. As such, it is likely that AF3 and Chai-1 are “overfitted” to the most common types of PLI structures in the PDB and may overlook several uncommon types of PLI binding pockets present in nature.
@@ 2.3 PoseBusters Benchmark results
With approximately half of its PLI structures deposited in the PDB after AF3’s maximum-possible training data cutoff of September 30, 2021 (n=308 total, filtered to n=130 for subsequent analyses), the PoseBusters Benchmark dataset’s results, presented in Figure 4, indicate once again that DL co-folding methods achieve top performance compared to conventional and DL docking baseline methods. Nonetheless, we observe an interesting phenomenon whereby Chai-1 strikes a balance of structural and chemical accuracy and chemical specificity comparable to AF3 even without input MSAs, potentially suggesting that Chai-1 achieves stronger sequence generalization for this dataset than AF3. Moreover, with the single-sequence version of AF3, we again observe significant degradations in its overall performance, whereas running Chai-1 with input MSAs achieves higher chemical specificity at the cost of marginal structural accuracy compared to running it in single-sequence mode. These observations highlight the importance in future work of carefully studying why and how the training of biomolecular structure generative models can be influenced to varying degrees by the availability and composition of diverse input MSAs.
@@ 2.4 CASP15 results
Figure 6: CASP15 single-ligand docking success rates (n=6).
As a new dataset of novel and challenging PLI complexes on which no method has been trained, the CASP15 dataset’s multi-ligand results (n=13), illustrated in Figure 5, indicate that most methods fail to adequately generalize to multi-ligand prediction targets, yet AF3 stands out in this regard (only) when provided input MSAs. As many of these CASP15 multi-ligand targets represent large, highly symmetric protein complexes, it is likely that additional evolutionary information in the form of MSAs has improved AF3’s ability to predict higher-order protein-protein interactions for these targets, yet interestingly its improved rate of structural accuracy comes at the cost of its protein-ligand chemical specificity (in comparison to its single-sequence results). For the CASP15 dataset’s single-ligand (i.e., primary ligand) results (n=6) presented in Figure 6, this trend is subverted in that conventional docking and simpler DL co-folding methods such as AutoDock Vina and NeuralPLexer outperform all recent DL co-folding methods in modeling crystalized PLIFs while achieving comparable rates of structural accuracy. Given the small size of the CASP15 dataset, it is reasonable to conclude that DL methods, in particular the latest co-folding methods, may be challenged to predict PLI complexes containing novel PLIs mediated by novel protein sequences. In the following Section 2.5, we will explore this latter point in greater detail by analyzing the protein sequence similarities between common PDB training data and this benchmark’s evaluation datasets.
Figure 7: Function annotations of the PLI complexes all methods mispredicted (n=129).
@@ 2.5 Exploratory analyses of results
Figure 8: Function annotations of the PLI complexes AF3 mispredicted (n=171).
@@
$\rightarrow$ To address this query, we first collect all ligand pose predictions that no method could predict with structural and chemical accuracy (according to the metric RMSD $\leq$ 2 Å & PB-Valid). For each of these “failed” ligand poses, we retrieve the PDB’s functional annotation of the protein in complex with this ligand and construct a histogram to visualize the frequency of these (failed complex) annotations. The results of this analysis are presented in Figure 7, in which we see that metal transport proteins, flavoproteins, biosynthetic proteins, RNA binding proteins, immune system proteins, and oxidoreductases are commonly mispredicted by all baseline methods such as Chai-1 and RoseTTAFold-All-Atom (RFAA) [7], suggesting these classes of proteins may be largely unaddressed by the most recent DL methods for PLI structure prediction. To illuminate potential future research directions, in the next analysis, we investigate whether this pattern persists specifically for one of the latest DL co-folding methods, AF3.
@@
Research Question 2: What are the most common types of protein-ligand complexes that DL co-folding methods such as AF3 fail to predict?
@@
$\rightarrow$ For this follow-up question, we link all of AF3’s failed ligand predictions with corresponding protein function annotations available in the PDB to understand which types of PLI complexes AF3 finds the most difficult to predict. Similar to the answer to our first research question, Figure 8 shows that, in order of difficulty, AF3 is most challenged to produce ligand poses of high structural and chemical accuracy for ligand-bound RNA binding proteins, immune system proteins, metal transport proteins, biosynthetic proteins, flavoproteins, lyases, and oxidoreductases. As several of these classes of proteins have not been well represented in the PDB over the last 50 years (e.g., immune system and biosynthetic proteins), in future work, it will be important to ensure that either the performance of new DL methods for PLI structure prediction is expanded to support accurate modeling of these uncommon types of ligand-bound proteins or a broadly applicable fine-tuning method for uncommon types of interactions is proposed.
Research Question 3: How often is lack of sequence homology to PDB training data associated with failed predictions by DL co-folding methods such as AF3?
$\rightarrow$ To understand the impact of protein sequence similarity on the performance of the DL co-folding method AF3, we isolate the subset of failed ligand pose predictions AF3 made for the PoseBusters Benchmark and CASP15 datasets, as none of these datasets’ prediction targets are contained in AF3’s training dataset. We then use MMseqs2 [21] to identify the deposition dates of the most similar (i.e., top hit) protein chains with 30% or greater sequence homology to any protein chain in the unseen PLI complexes AF3 failed to predict. Figure 9 reveals that most of the unseen PLI complexes AF3 failed to predict were not associated with any protein sequence homologs present in its PDB-based training dataset. That is, when AF3 failed to predict a new PLI complex, it also could not rely on sequence homology to its training dataset to bolster its performance. This observation suggests that the performance of recent DL co-folding methods for novel protein sequences or PLI complexes may be limited by the extent to which the method can ”retrieve” similar sequence representations from its training data. We conclude our quantitative analyses with an illustration of the different failure modes of each baseline method, as depicted in Figure 10.
Figure 9: Sequence homologs of the unseen PLI complexes AF3 mispredicted (n=62).
@@
Figure 10: Examples of baseline methods’ three failure modes discussed in this work.

### v6

# Assessing the potential of deep learning for protein-ligand docking

@@ Abstract
The effects of ligand binding on protein structures and their in vivo functions carry numerous implications for modern biomedical research and biotechnology development efforts such as drug discovery. Although several deep learning (DL) methods and benchmarks designed for protein-ligand docking have recently been introduced, to date no prior works have systematically studied the behavior of the latest docking and structure prediction methods within the broadly applicable context of (1) using predicted (apo) protein structures for docking (e.g., for applicability to new proteins); (2) binding multiple (cofactor) ligands concurrently to a given target protein (e.g., for enzyme design); and (3) having no prior knowledge of binding pockets (e.g., for generalization to unknown pockets). To enable a deeper understanding of docking methods’ real-world utility, we introduce PoseBench, the first comprehensive benchmark for broadly applicable protein-ligand docking. PoseBench enables researchers to rigorously and systematically evaluate DL methods for apo-to-holo protein-ligand docking and protein-ligand structure prediction using both primary ligand and multi-ligand benchmark datasets, the latter of which we introduce for the first time to the DL community. Empirically, using PoseBench, we find that (1) DL co-folding methods generally outperform comparable conventional and DL docking baseline algorithms, yet popular methods such as AlphaFold 3 are still challenged by prediction targets with novel protein-ligand binding poses; (2) certain DL co-folding methods are highly sensitive to their input multiple sequence alignments, while others are not; and (3) DL methods struggle to strike a balance between structural accuracy and chemical specificity when predicting novel or multi-ligand protein targets. Code, data, tutorials, and benchmark results are available at https://github.com/BioinfoMachineLearning/PoseBench.
@@ 1 Introduction
As the field has gradually begun to investigate whether proteins in complex with other types of molecules can faithfully be modeled with ML (and particularly deep learning (DL)) techniques [5, 6, 7], several new works in this direction have suggested the promising potential of such approaches to protein-ligand structure determination [8, 9, 10, 11]. Nonetheless, it remains to be shown the extent to which the latest of such (docking and co-folding-based) DL methods can adequately generalize to the context of binding novel or uncommon protein-ligand interaction (PLI) pockets and multiple interacting ligand molecules (e.g., which can alter the chemical functions of various enzymes) as well as whether such methods can faithfully model amino acid-specific types of PLIs natively found in crystallized biomolecular structures.
Figure 1: Overview of PoseBench, our comprehensive benchmark for broadly applicable DL modeling of primary and multi-ligand protein complex structures. Baseline algorithms within the benchmark include a range of the latest DL docking and co-folding methods, both open-source and commercially restrictive, as well as conventional algorithms for docking. Key observations derived using PoseBench include the discontinuity between structure and interaction modeling performance for novel or uncommon prediction targets and the heavy reliance of key DL co-folding methods on MSA-based input features to achieve high structural accuracy.
@@
  • We introduce the first unified benchmark for protein-ligand docking and structure prediction that evaluates the performance of several recent DL-based baseline methods (DiffDock-L, DynamicBind, NeuralPLexer, RoseTTAFold-All-Atom, Chai-1, Boltz-1, and AlphaFold 3) as well as conventional algorithms (P2Rank + AutoDock Vina) for primary and multi-ligand docking, which suggests that DL co-folding methods generally outperform conventional algorithms yet remain challenged by novel or uncommon prediction targets.
  • In contrast to several recent works using crystal protein structures for protein-ligand docking [12, 13], the docking benchmark results we present in this work are all within the context of standardized input multiple sequence alignments (MSAs) and high accuracy apo-like (i.e., AlphaFold 3-predicted) protein structures (see Appendix D) without specifying known binding pockets, which notably enhances the broad applicability of this study’s findings.
  • Our newly proposed benchmark, PoseBench, enables specific insights into necessary areas of future work for accurate and generalizable biomolecular structure prediction, including that DL methods struggle to balance faithful modeling of native PLI fingerprints (PLIFs) with structural accuracy during pose prediction and that some DL co-folding methods (AlphaFold 3) are more dependent than others (Boltz-1, Chai-1) on the availability of input MSAs.
  • Our benchmark results also highlight the importance of including challenging (out-of-distribution) datasets when evaluating future DL methods while measuring their ability to recapitulate amino acid-specific PLIFs with an appropriate new metric that we introduce in this work.
@@ 2.1 Astex Diverse results
Containing PLI structures deposited in the RCSB Protein Data Bank (PDB) [15] up until 2007, most of the well-known Astex Diverse dataset’s structures [16] are present in the training data of each baseline method, yet benchmarking results for this dataset (n=85), shown in Figure 2, indicate that only DL co-folding methods achieve higher structural and chemical accuracy rates (RMSD $\leq$ 2 Å & PB-Valid) than the conventional docking baseline AutoDock Vina combined with P2Rank for PLI binding site prediction to facilitate blind molecular docking. Interestingly, nearly all baseline methods identify the correct PLI binding pocket approximately 90% of the time, yet only the DL co-folding methods Chai-1 [17], Boltz-1 [18], and AlphaFold 3 (AF3) [11] achieve a reasonable balance between their rates of structural and chemical accuracy and chemical specificity (PLIF-WM), with the single-sequence (i.e., MSA-ablated) version of AF3 being a notable exception. These results suggest that DL co-folding methods have learned the most comprehensive representations of this dataset’s input sequences, yet only the DL co-folding method Chai-1 maintains strong performance without the availability of diverse input MSAs. One likely explanation for this phenomenon is that Chai-1’s training relied upon the availability of amino acid sequence embeddings generated by the protein language model ESM2 [19] in addition to features derived from input MSAs, which may have imbued the model with rich MSA-independent representations for biomolecular structure prediction.
@@ 2.2 DockGen-E results
As visualized in Figure 3, results with our new DockGen-E dataset of biologically relevant PLI complexes deposited in the PDB up to 2019 (n=122) demonstrate that only the latest DL co-folding methods can locate a sizable fraction of structurally accurate PLI binding poses represented in this dataset. As such methods may have previously seen these PLI structures in their respective training data, it is surprising that even the latest AF3 model fails to identify a structurally and chemically accurate pose for more than 75% of the dataset’s complexes. Further, for Chai-1, Boltz-1, and AF3, their single-sequence variants achieve higher chemical specificity than their MSA-based versions, which may indicate that for these methods MSA features obfuscate primary sequence knowledge in favor of evolution-averaged (i.e., amino acid-generic) representations. The overall lower range of PLIF-WM values achieved by each method for this dataset further suggests the increased chemical modeling difficulty of this dataset’s complexes compared to those presented by the Astex Diverse dataset. A potential source of these difficulties is that each of this dataset’s complexes represents a functionally distinct PLI binding pocket (as codified by ECOD domains [20], see [21] for more details) compared to data deposited in the PDB before 2019. As such, it is likely that Chai-1, Boltz-1, and AF3 are “overfitted” to the most common types of PLI structures in the PDB and may overlook several uncommon types of PLI binding pockets present in nature.
@@ 2.3 PoseBusters Benchmark results
With approximately half of its PLI structures deposited in the PDB after AF3 and Boltz-1’s maximum-possible training data cutoff of September 30, 2021 (n=308 total, filtered to n=130 for subsequent analyses), the PoseBusters Benchmark dataset’s results, presented in Figure 4, indicate once again that DL co-folding methods achieve top performance compared to conventional and DL docking baseline methods. Nonetheless, we observe an interesting phenomenon whereby Chai-1 strikes a balance of structural and chemical accuracy and chemical specificity comparable to that of the best-performing AF3, even without input MSAs, potentially suggesting that Chai-1 achieves stronger binding pose generalization for this dataset than AF3. Moreover, with the single-sequence version of AF3, we again observe significant degradations in its overall performance, whereas running Chai-1 with input MSAs achieves higher chemical specificity at the cost of marginal structural accuracy compared to running it in single-sequence mode. These observations highlight the importance in future work of carefully studying why and how the training of biomolecular structure generative models can be influenced to varying degrees by the availability and composition of diverse input MSAs.
@@ 2.4 CASP15 results
Figure 6: Extended Data Figure 1 — CASP15 single-ligand docking success rates (n=6).
As a new dataset of novel and challenging PLI complexes on which no method has been trained, the CASP15 dataset’s multi-ligand results (n=13), illustrated in Figure 5, indicate that most methods fail to adequately generalize to multi-ligand prediction targets, yet AF3 stands out in this regard (only) when provided input MSAs. As many of these CASP15 multi-ligand targets represent large, highly symmetric protein complexes, it is likely that additional evolutionary information in the form of MSAs has improved AF3’s ability to predict higher-order protein-protein interactions for these targets, yet interestingly its improved rate of structural accuracy comes at the cost of its protein-ligand chemical specificity (in comparison to its single-sequence results). For the CASP15 dataset’s single-ligand (i.e., primary ligand) results (n=6) presented in Figure 6, this trend is subverted in that conventional docking and open-source DL co-folding methods such as AutoDock Vina, NeuralPLexer, and Boltz-1 outperform all other recent DL co-folding methods in modeling crystalized PLIFs while achieving comparable rates of structural accuracy. Given the small size of the CASP15 dataset, it is reasonable to conclude that DL methods, in particular some of the latest co-folding methods, may be challenged to predict protein-ligand complexes containing novel PLIs. In the following Section 2.5, we will explore this latter point in greater detail by analyzing the protein-ligand binding similarities between common PDB training data and this benchmark’s evaluation datasets.
Figure 7: Extended Data Figure 2 — Function annotations of the PLI complexes all methods mispredicted (n=129).
@@ 2.5 Exploratory analyses of results
Figure 8: Extended Data Figure 3 — Function annotations of the PLI complexes AF3 mispredicted (n=171).
@@
$\rightarrow$ To address this query, we first collect all ligand pose predictions that no method could predict with structural and chemical accuracy (according to the metric RMSD $\leq$ 2 Å & PB-Valid). For each of these “failed” ligand poses, we retrieve the PDB’s functional annotation of the protein in complex with this ligand and construct a histogram to visualize the frequency of these (failed complex) annotations. The results of this analysis are presented in Figure 7, in which we see that metal transport proteins, flavoproteins, biosynthetic proteins, RNA binding proteins, immune system proteins, and oxidoreductases are commonly mispredicted by all baseline methods such as Chai-1 and RoseTTAFold-All-Atom (RFAA) [7], suggesting these classes of proteins may be largely unaddressed by the most recent DL methods for PLI structure prediction. To illuminate potential future research directions, in the next analysis, we investigate whether this pattern persists specifically for AF3, the most accurate DL co-folding method according to our benchmarking results.
@@
Research Question 2: What are the most common types of protein-ligand complexes that highly-accurate DL co-folding methods such as AF3 fail to predict?
@@
$\rightarrow$ For this follow-up question, we link all of AF3’s failed ligand predictions with corresponding protein function annotations available in the PDB to understand which types of PLI complexes AF3 finds the most difficult to predict (to understand its predictive coverage of important molecular functions). Similar to the answer to our first research question, Figure 8 shows that, in order of difficulty, AF3 is most challenged to produce ligand poses of high structural and chemical accuracy for ligand-bound RNA binding proteins, immune system proteins, metal transport proteins, biosynthetic proteins, flavoproteins, lyases, and oxidoreductases. As several of these classes of proteins have not been well represented in the PDB over the last 50 years (e.g., immune system and biosynthetic proteins), in future work, it will be important to ensure that either the performance of new DL methods for PLI structure prediction is expanded to support accurate modeling of these uncommon types of ligand-bound proteins or a broadly applicable fine-tuning method for uncommon types of interactions is proposed.
Research Question 3: Is lack of protein-ligand binding pose homology to PDB training data (inversely) correlated with the prediction accuracy of each method?
Figure 9: Extended Data Figure 4 — Relationship between methods’ PoseBusters Benchmark prediction accuracy (RMSD) and each complex’s training set similarity (SuCOS-pocket similarity) (n=41).
@@
Figure 10: Extended Data Figure 5 — Examples of baseline methods’ three failure modes discussed in this work.
$\rightarrow$ To understand the impact of protein-ligand binding pose similarity on the performance of each baseline method, we use the PLINDER data resource [22] to identify (n=41/130) cluster representatives of the PoseBusters Benchmark dataset based on the product of each complex’s ligand (structure and feature-based) Combined Overlap Score (SuCOS) [23] and its protein binding pocket’s (structure and sequence-based) similarity [24], as none of this subset’s prediction targets are contained in any method’s training dataset. For these cluster representatives, we then calculate the Pearson and Spearman correlations (and p-values) between each method’s complex prediction accuracy (i.e., ligand pose RMSD) and the complex’s maximum SuCOS-binding pocket-based similarity to any complex deposited in the PDB before AF3’s training dataset cutoff of September 30, 2021. Figure 9 reveals that all DL methods’ performance is correlated with a complex’s similarity to common PDB training data, with (MSA-based) Boltz-1, AF3, and Chai-1 exhibiting the strongest and most statistically significant ($p<0.05$) correlations. Like concurrent work assessing the performance of co-folding methods in novel prediction settings [24], our findings suggest that although the current generation of deep learning models (both docking and co-folding methods) for protein-ligand docking and structure prediction can occasionally make accurate predictions for truly novel (SuCOS-pocket similarity << 30) protein-ligand complexes, such methods rely (at least in large part) on recapitulating protein-ligand binding patterns seen during training to make accurate predictions for unseen complexes (n.b., while interestingly for conventional docking methods such as P2Rank-Vina, this trend is not observed). We conclude our quantitative analyses with an illustration of the different failure modes of each baseline method, as depicted in Figure 10. In this figure, we illustrate that DL methods such as RFAA and AF3 commonly struggle to accurately predict the structure of ligand-binding biosynthetic and immune system proteins, suggesting that these (uncommon) types of PLIs are not well addressed by the current generation of DL-based structure prediction methods, suggesting future research opportunities for interaction-specific modeling (e.g., through the use of fine-tuning or preference optimization).